# JANUS : A Difference-Oriented Analyzer For Financial Centralization Risks in Smart Contracts

### v1

@@ VI Limitations
Focusing on financial centralization risks. JANUS focuses on detecting centralization risks that affect financial variables, such as token balances, as these can directly cause financial losses to contract users. While our current implementation cannot automatically detect risks that do not affect financial variables (which we did not encounter in our experiments), JANUS offers flexibility through a customizable interface. This interface allows users to specify target variables, enabling the detection of non-financial centralization risks without using our financial variable recognition module. Additionally, for cases where JANUS does not accurately identify financial variables, users can manually input the variable names.
Implementation based on another tool. Some key steps in JANUS, e.g., symbolic execution and data flow analysis, are implemented based on an existing tool Slither. As a result, we cannot definitively determine if the symbolic execution process generates infeasible states or if the data flow analysis leads to omissions in functions that require analysis or variables that need labeling. However, our comparative analysis using the comparison dataset does not reveal any such issues, suggesting the reliability of our tool. For our real-world dataset consisting of 33,151 unlabeled contracts, we do not perform a comprehensive manual check for false positive results which would be labor-intensive.
@@ VII Conclusion
We propose and implement JANUS, which can detect financial centralization risks independently of their behavior patterns. JANUS employs a difference-oriented state traversal method and variable summaries to reduce the number of states to be compared. We prove that these methods do not introduce false alarms or cause omissions in centralization risk detection. JANUS outperforms other automatic tools in terms of risk detection accuracy, and it successfully finds two types of centralization risks that other automatic tools fail to detect.
Appendix A The Iterative Algorithm for Obtaining Differences in Labeled States
In this algorithm, $labeled\_symbolic\_exec(f,a,l)$ symbolically executes the function $f$ using account $a$, and label the variables dependent on the labeled variables in $l$.
@@ Appendix A The Iterative Algorithm for Obtaining Differences in Labeled States
1:$F\leftarrow$ the functions in $C$
2:$l_{0}\leftarrow$ initial labeled state after executing the constructor of $C$ (The set of labeled variables is $\emptyset$)
3:$D\leftarrow\emptyset$
4:$S_{next}\leftarrow\emptyset$
@@
Appendix B The Experiments about the Hyperparameters of Our Financial Variable Recognition Module
@@ Appendix C Theorems and Proofs
TABLE V: The Design of Variable Summaries for Different Variable Types
@@
1) For all $v\in(BV(s_{i+1},s_{i+1}^{\prime})\setminus BV(s_{i},s_{i}^{\prime}))$, $v$ is dependent of variables in $BV(s_{i},s_{i}^{\prime})$.
@@
By Definition of exec_state and TABLE V, the value of exec_state is corresponding to the values of keys in its summary. Therefore, if $l_{n}.\sigma(v)\neq l_{n}^{\prime}.\sigma(v)$, we have $\phi(l_{n-1},l_{n},v)\neq\phi(l_{n-1},l_{n}^{\prime},v)$. Theorem 3 holds.
@@
2) By TABLE V, the keys of $\phi(l,l_{n},v)$ is $K_{p}$ and the keys of $\phi(l,l_{n}^{\prime},v)$ is $K_{o}$.

### v2

@@ VI Limitations and Future Work
Limited Detection Scope. Our tool currently focuses on detecting centralized risks primarily associated with finance-related variables (e.g., token balances) and privileged accounts identified from state variables of the address type. Consequently, our tool may miss non-financial risks (e.g., mutable metadata [48]) or privileged accounts represented using other types, potentially resulting in false negatives. As a supplement, we provide a customizable interface that allows users to specify target variables manually, thereby extending detection to non-financial risks. It is noteworthy that even within this limited scope, our tool has successfully identified practical centralized risk cases overlooked by existing tools, demonstrating its effectiveness and value. In the future, we plan to integrate large language models, which have been used by multiple studies[38][49] to understand smart contract semantics, to identify a broader range of key variables and privileged account variables.
Dependency on External Static Analysis Engine. JANUS leverages Slither for foundational static analysis to avoid reinventing established components. As a result, we cannot definitively determine if the symbolic execution process generates infeasible states or if the data flow analysis leads to omissions in functions that require analysis or variables that need labeling. However, it is important to note that our iterative algorithm is independent and operates on an intermediate representation. This means that while Slither is currently used, our method could be adapted to other analysis engines (e.g., Mythril) that provide a similar code representation. The current implementation’s reliability is supported by its performance on the comparison dataset, where no obvious issues from Slither’s analysis are observed. In the future, we plan to integrate other static analysis tools and employ cross-validation to enhance the reliability of foundational analysis and develop a generic interface to enable more flexible integration of our core algorithms with different static analysis backends.
@@
The trade-off between accuracy and efficiency. While our tool’s execution speed is currently slower than that of pattern-based matching tools, limiting its practical deployment, it offers superior detection accuracy and the capability to identify unknown risks through symbolic execution. These advantages are critical in real-world applications. In the future, we plan to further optimize the tool’s efficiency through methods such as parallel analysis of multiple contracts and incorporating preliminary screening with other tools.
@@ VII Conclusion
We propose and implement JANUS, which can detect financial centralized risks independently of their behavior patterns. JANUS employs a difference-oriented state traversal method and variable summaries to reduce the number of states to be compared. We prove that these methods do not introduce false alarms or omissions in detection. JANUS outperforms other tools in terms of risk detection accuracy, and it successfully finds two types of risks that other tools fail to detect.
Appendix A Appendix
A-A The Iterative Algorithm for Obtaining Differences in Labeled States
To facilitate understanding, a simplified state representation is used in the algorithm in our manuscript. The algorithm we employ in our tool is shown below, which follows the same main flow but replaces the states with labeled states. In this algorithm, labeled_symbolic_exec(f,a,l)labeled\_symbolic\_exec(f,a,l) symbolically executes the function $f$ using account $a$, and label the variables dependent on the labeled variables in $l$.
@@ A-A The Iterative Algorithm for Obtaining Differences in Labeled States
[…]
@@
A-B The Experiments about the Hyperparameters of Our Financial Variable Recognition Module
@@ A-C Theorems and Proofs
TABLE IV: The Design of Variable Summaries for Different Variable Types
@@
1) For all $\in(BV(s_{i+1},s_{i+1}^{\prime})\setminus BV(s_{i},s_{i}^{\prime})), vv$ is dependent of variables in $BV(s_{i},s_{i}^{\prime})$.
@@
By Definition of exec_state and TABLE IV, the value of exec_state is corresponding to the values of keys in its summary. Therefore, if $l_{n}.\sigma(v)\neq l_{n}^{\prime}.\sigma(v)$, we have $\phi(l_{n-1},l_{n},v)\neq\phi(l_{n-1},l_{n}^{\prime},v)$. Theorem 3 holds.
@@
2) By TABLE IV, the keys of $\phi(l,l_{n},v)$ is $K_{p}$ and the keys of $\phi(l,l_{n}^{\prime},v)$ is $K_{o}$.